# Clickbait Detection via Large Language Models

### v1

@@ 4.3 Implementation details and evaluation metrics
Paper CNN 68.87 73.27 68.76 66.11 71.45 74.28 70.47 71.16 72.38 72.12 71.32 72.29
BiLSTM 64.36 68.41 65.96 65.89 68.44 68.07 66.26 64.94 69.78 68.23 69.03 69.85
GATED_CNN 68.44 72.40 68.63 67.29 71.41 74.78 69.36 65.84 71.79 72.02 71.04 72.91
FAST_TEXT 67.93 70.25 66.03 65.53 68.45 72.88 68.96 67.85 71.10 72.15 72.23 70.39
BERT 50.16 50.20 40.02 44.54 51.92 52.21 41.51 46.26 54.17 54.48 43.60 48.44
ChatGPT 68.33 72.18 62.70 66.42 68.32 70.40 63.88 66.98 70.94 72.14 67.21 69.59
Implementation details.
We chose the latest available LLMs GPT3.5 (gpt-3.5-turbo). In order to interact with the model, we access it by making API calls to the ChatGPT API, in which calling GPT3.5 requires payment. This model’s maximum context length is 4096 tokens, approximating 100 news headlines. Moreover, this model can only call three news headlines a minute. In addition, for the few-shot experiments, we select a few sentences that ChatGPT got wrong from the sevendata sets mentioned above as examples of clickbait detection. Figure 1 shows ChatGPT detecting clickbait errors.
@@
Table 7: Comparison of different prompts for GPT3.5 on English datasets (%).
@@
Table 8: Comparison of different prompts for GPT3.5 on Chinses datasets (%).
Data Method Shot
5 10 20
@@
Evaluation metrics
$Acc=\frac{TP+TN}{TP+TN+FP+FN}$
$P=\frac{TP}{TP+FP}$
$R=\frac{TP}{TP+FN}$
$F1=\frac{{2}\times{P}\times{R}}{P+R}$
4.4 Experimental Results
The main results of the experiments on clickbait detection in English and Chinese datasets on zero-shot and few-shot scenarios are listed in Table 4, 5 and 6 respectively. We can see that the performance of ChatGPT has not demonstrated the best results on all four metrics over the seven datasets. Compared to the deep neural networks and fine-tuning PLMs, ChatGPT has a lot of room for improvement in clickbait detection. Moreover, with the increasing of pre-trained information, the effectiveness of clickbait detection, sometimes, is worse. Therefore, we can observe that a small amount of pre-training information does not significantly impact the performance of ChatGPT in detecting clickbait. Moreover, we can observe that ChatGPT can achieve stable results on almost all evaluation metrics. The results confirmed that LLMs can be adapted to other languages, and ChatGPT is a monolithic model capable of supporting multiple languages, which makes it a comprehensive multilingual clickbait detection technique.
4.5 Ablation Study
We compared the results between different prompts (as shown in Table 1 and Table 2) for clickbait detection tasks. As same as the main experiments, we select GPT3.5 as a backbone for the prompts experiments. The results comparison of different Prompts is shown in Table 7 and Table 8.
Generally speaking, there is no significant performance gap between the first prompt (P1) and the second prompt (P2) in terms of effectiveness for few-shot clickbait detection. Specifically, the performance of P1 is significantly better than that of P2 in the dataset Wechat. In dataset Sina, the opposite is true.
@@ 5 Conclusion and Future Work
In this paper, we present a study of the performance of LLMs (ChatGPT with GPT3.5) for clickbait detection. During the benchmark experiments on both English and Chinese datasets, LLMs did not perform as well as the current state-of-the-art deep neural networks and fine-tuning PLMs methods in the realm of multilingual clickbait detection. In our subsequent efforts, we will try to design more effective methods with the help of LLMs that can significantly and consistently outperform SOTA clickbait detection methods.

### v4

@@ 4.3.3 Evaluation metrics
Paper CNN 72.21 75.57 72.21 73.85 76.20 74.09 76.20 75.13 79.33 77.48 79.33 78.41
BiLSTM 74.69 78.10 74.69 76.35 77.28 79.06 77.28 78.16 80.40 80.36 80.40 80.38
GATED-CNN 68.44 72.40 68.63 67.29 71.41 74.78 69.36 65.84 71.79 72.02 71.04 72.91
FAST-TEXT 67.93 70.25 66.03 65.53 68.45 72.88 68.96 67.85 71.10 72.15 72.23 70.39
BERT 75.79 76.99 75.79 76.38 79.25 77.63 79.25 78.42 82.65 82.71 82.65 82.68
MFWCD-BERT 78.75 80.01 78.69 78.90 80.84 80.83 80.79 81.03 84.54 85.02 84.11 84.51
PT 72.07 76.98 72.24 71.33 75.61 76.05 75.61 75.51 80.41 80.55 80.41 80.39
PEPL 70.14 74.46 70.14 70.53 73.62 73.88 73.62 73.55 78.28 78.77 78.28 78.19
[…]
GPT-4 77.75 80.04 73.85 76.82 72.6 66.99 89.09 76.48 62.28 57.34 95.82 71.75
4.4 Experimental Results
The main results of the experiments on clickbait detection in English and Chinese datasets are listed in Tables 3 and 4 respectively, which presents the detection results of ChatGPT (including GPT-3.5 and GPT-4) in the few-shot scenario, along with a comparison to other state-of-the-art methods. We can see that the performance of ChatGPT has not demonstrated the best results on all four metrics over the seven datasets. Compared to the deep neural networks and fine-tuning PLMs, ChatGPT has a lot of room for improvement in clickbait detection, especially the performance of GPT-3.5 on the Chinese dataset. In addition, with the increase of pre-trained information, the effectiveness of clickbait detection, sometimes, is worse. Therefore, we can observe that a small amount of pre-training information does not significantly impact the performance of ChatGPT in detecting clickbait. Moreover, we can observe that ChatGPT can achieve stable results on almost all evaluation metrics. The results confirmed that LLMs can be adapted to other languages, and ChatGPT is a monolithic model capable of supporting multiple languages, which makes it a comprehensive multilingual clickbait detection technique. It is worth noting that GPT-4 shows significant improvements over GPT-3.5 in areas such as comprehension and handling complex tasks, especially in scenarios that require a high level of text understanding, such as clickbait detection tasks.
Moreover, we have provided a detailed analysis of four news headlines where ChatGPT correctly classified them using the prompt method but made errors when detecting clickbait. For example, given the English not-clickbait headline ”Tigers ’starved to death’ to make \$500 aphrodisiac wine with their bones,” ChatGPT misclassified it as clickbait. This is primarily because, first, the headline touches on sensitive topics such as animal abuse, illegal wildlife trade, and ethical controversies, which can easily elicit emotional responses from readers. Secondly, terms like ”starved to death” and ”aphrodisiac” are strong verbs and adjectives often used exaggeratedly or emotionally, and these words may match common vocabulary found in clickbait articles. Thirdly, in the absence of additional context, ChatGPT may be more inclined to categorize sentences with controversial or extreme characteristics as clickbait.For the Chinese not-clickbait headline ”扬州通报“网传副市长与副局长生活作风问题”:建议对二人免职” (Yangzhou reports ’alleged misconduct in the personal conduct of deputy mayor and deputy director’: Suggests dismissal of the two individuals.), ChatGPT misclassified it as clickbait. This is mainly because it is an attention-grabbing topic with relatively complex grammatical structures, including multiple parallel phrases and verb phrases. This complexity can lead the model to struggle in interpreting the sentence. It’s also important to note that the phrase ”网传” (spread on the internet) may have contributed to ChatGPT’s classification as spreading false information.For the English clickbait headline ”This is America’s favorite fast food restaurant,” ChatGPT misclassified it as not-clickbait. This is primarily because it is a topic related to American food culture and consumption habits, falling under the category of everyday life. Furthermore, the headline uses proper grammatical structure to pose a clear question, making it appear as a reasonable query rather than clickbait.For the Chinese clickbait headline ”与你有关!一批重要国家标准,今天发布!” (Related to You! A batch of important national standards released today!), ChatGPT misclassified it as not-clickbait. This is mainly because national standards and regulations are a topic of practical significance, and the proper grammatical structure is used to present a statement. These factors may have led ChatGPT to mistakenly classify it as a legitimate news headline rather than clickbait.
In general, ChatGPT performs well in handling objective questions with clear, standardized answers. However, when it comes to addressing personalized questions, which often lack a fixed standard answer, it can be influenced by certain specific features. Moreover, conducting multiple tests on the same news may yield different results, leading to suboptimal performance in metrics like accuracy for clickbait detection.
4.5 Ablation Study
We compared the results between different prompts (as shown in Table 12) for clickbait detection tasks. As same as the main experiments, we select GPT-3.5 as a backbone for the prompts experiments. The results comparison of different Prompts are shown in Table 5.
Generally speaking, there is no significant performance gap between the first prompt (P1) and the second prompt (P2) in terms of effectiveness for clickbait detection. Specifically, the performance of P1 is significantly better than that of P2 in the dataset WeChat. In dataset Sina, the opposite is true.
@@ 4.5 Ablation Study
Table 5: Comparison of different prompts for GPT-3.5(%). The bolder ones mean better.
@@
Comparison of different prompts for GPT-3.5 on Chinese datasets
@@
4.6 The Impact of Different LLMs
Due to the rapid development and diverse characteristics of LLMs, we selected several different LLMs as the foundation for our experiments on clickbait detection. Specifically, we chose local LLMs including LLaMA3-8B, Yi-34B, and DeepSeek-32B, and compared their performance with commercial models such as GPT-3.5 and GPT-4. The results are presented in Tables 6 and 7.
We observe that GPT-4 consistently performs well across most datasets. Among the local LLMs, DeepSeek-32B demonstrates relatively stable performance and even surpasses GPT-3.5 in certain metrics. However, it is important to note that some LLMs tend to favor specific classes, resulting in a significant imbalance between precision and recall. As a result, LLMs still show considerable limitations in the task of clickbait detection.
Table 6: The Accuracy, Precision, Recall, F1-scores results (%) on English datasets with different LLMs(%). The bolder ones mean better.
DLC CND SCC
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
LLaMA3 77.68 66.41 69.67 68.00 59.32 61.52 49.77 55.02 82.95 88.78 75.43 81.56
Yi 76.71 82.91 39.75 53.74 58.19 58.81 54.66 56.66 79.10 84.12 70.58 76.92
DeepSeek 72.11 57.80 66.80 61.98 62.85 62.53 64.13 63.32 84.13 90.22 76.57 82.83
GPT-3.5 80.75 69.06 82.79 73.56 60.00 59.02 68.36 62.06 83.94 88.63 90.01 83.08
GPT-4 90.54 94.00 77.05 84.68 68.61 69.47 66.97 68.20 94.62 98.83 90.30 94.37
Table 7: The Accuracy, Precision, Recall, F1-scores results (%) on Chinese datasets with different LLMs(%). The bolder ones mean better.
Sina Tencent WeChat Paper
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
LLaMA3 73.52 69.12 85.02 76.25 63.87 60.04 82.98 69.67 54.64 52.91 84.52 65.08 70.63 64.42 92.16 75.83
Yi 68.89 71.07 69.97 70.55 72.54 68.87 73.46 71.08 63.68 61.02 75.71 67.52 76.45 76.54 76.30 76.30
DeepSeek 82.82 93.67 76.83 84.42 78.45 77.35 80.47 79.03 74.66 71.01 83.33 76.66 79.39 88.32 74.70 80.89
GPT-3.5 70.73 76.21 60.28 67.32 59.34 57.91 71.13 63.84 60.12 56.69 85.71 68.24 68.33 72.18 62.70 66.42
GPT-4 73.87 93.08 75.57 76.96 65.97 67.76 75.94 71.58 73.79 69.18 88.57 77.68 77.75 80.04 73.85 76.82
@@ 5 Conclusion and Future Work
In this paper, we present a study of the performance of LLMs (ChatGPT with GPT-3.5 and GPT-4) for clickbait detection. During the benchmark experiments on both English and Chinese datasets, LLMs did not perform as well as the current state-of-the-art deep neural networks and fine-tuning PLMs methods in the realm of multilingual clickbait detection. In our subsequent efforts, we will try to design more effective methods with the help of LLMs that can significantly and consistently outperform SOTA clickbait detection methods.